February 3, 2004
Linda Bartholomee, Mess Attendant, makes sure the crew is well fed and she always provides a healthy serving of smiles.
Most of the science party awoke to find the Atlantis was no longer moving forward. We had finally arrived at 1 pm Tuesday at our first area for Alvin dives and camera tows. Almost immediately the camera tow prepared for a traverse that covered terrain up to 3 km east of the EPR axis. The “tow team” pulled its first “all nighter” with Dan flying the "Rabbit" cam tow, and Rhian and Javier navigating. The camera system worked beautifully as did the magnetometer, so we now have our first set of bottom images (see below) and first data on variations in the magnetic field along the tow track. We triggered 8 waxed corer balls but lost three to the lava flows. They must have become wedged between the lobate and pillowed flows, snapping the retrieval lines. The 5 we did retrieve all collected volcanic glass. So, although the score (Fornari -5, EPR Lava Flows-3) gave us an away win, we continue to tweak the system hoping for a series of shutouts.
The data screen from the Rabbit Cam showing the trace from the new forward looking, obstacle avoidance altimeter (yellow line), and a sucessful wax "bomb" with glass samples.
Three images of the volcanic seafloor on the east side of the EPR axis near 9° 53'N collected during Camera Tow#1 last night. The middle photo shows one of the wax balls being recovered by the small winches on the TowCam. Right photo shows a pink crinoid attached to the pillow lava.
Alvin was lowered into the water a little after 8 am and sank beneath the surface about 8:30 am. This followed a hectic night for the Alvin crew, who had to swap out a malfunctioning manipulator arm, and reflects the continuous monitoring of Alvin functions prior to a dive and dedication of the team to getting science done at the seafloor.
Left- Tony Tarantino, one of the Alvin Pilots looks at the full load in the Alvin basket for today's dive. Right- Bill Seyfried framed by Alvin's port manipulator, the one that was serviced last night, as Alvin is pulled out of the hanger for today's dive.
Today was a biology and geochemistry dive with Kang Ding and Stacey Beaulieu deploying several time-series experiments at high and low temperature hydrothermal vents. Remote monitoring of the geochemistry probes the U. Minnesota group deployed from Alvin indicates that all are working well.
Canjun Yang preparing a geochemistry probe for the February 3rd Alvin dive
Atlantis waved goodbye to the "Rat" Cam... for now.
Alvin will 'pose' for the camera and retreive it on an upcoming dive
HAPPY BIRTHDAY ADAM